2. Consider a linear transformation T: R3 → R3 defined by



  ([x         [ x + 4y +3z  

T  y     =      -5y - 4z 

    z])        5x + 10y + 7z ]



Note: T is a 3x1 matrix containing x, y, z respectively. T is equal to another matrix as shown above.


a) Find the matrix A for T


b) Find a basis for ker(T) and the dim(ker(T)). Then find dim(Im(T)), without finding a basis for Im(T). (Show all working)


c) Find a basis for Im(T)



1
Expert's answer
2022-03-29T01:14:53-0400

"a:The\\,\\,matrix\\,\\,is\\,\\,of\\,\\,coefficients\\,\\,written\\,\\,in\\,\\,lines:\\\\A=\\left[ \\begin{matrix}\t1&\t\t4&\t\t3\\\\\t0&\t\t-5&\t\t-4\\\\\t5&\t\t10&\t\t7\\\\\\end{matrix} \\right] \\\\b:\\\\A\\left[ \\begin{array}{c}\tx\\\\\ty\\\\\tz\\\\\\end{array} \\right] =0\\Rightarrow \\left\\{ \\begin{array}{c}\tx+4y+3z=0\\\\\t-5y-4z=0\\\\\t5x+10y+7z=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\tx+4y-\\frac{15}{4}y=0\\\\\t5x+10y-\\frac{35}{4}y=0\\\\\tz=-\\frac{5}{4}y\\\\\\end{array} \\right. \\Rightarrow \\\\\\Rightarrow \\left\\{ \\begin{array}{c}\ty=-4x\\\\\ty=-4x\\\\\tz=-\\frac{5}{4}y\\\\\\end{array} \\right. \\Rightarrow \\left[ \\begin{array}{c}\tx\\\\\ty\\\\\tz\\\\\\end{array} \\right] =t\\left[ \\begin{array}{c}\t1\\\\\t-4\\\\\t5\\\\\\end{array} \\right] \\,\\,-\\,\\,one\\,\\,vector\\\\dim\\left( ker\\left( T \\right) \\right) =1\\\\dim\\left( \\mathrm{Im}\\left( T \\right) \\right) =3-dim\\left( ker\\left( T \\right) \\right) =3-1=2\\\\c:\\\\A\\left[ \\begin{array}{c}\t1\\\\\t0\\\\\t0\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t1\\\\\t0\\\\\t5\\\\\\end{array} \\right] ,A\\left[ \\begin{array}{c}\t0\\\\\t1\\\\\t0\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t4\\\\\t-5\\\\\t10\\\\\\end{array} \\right] \\\\Two\\,\\,vectors\\,\\,\\left[ \\begin{array}{c}\t1\\\\\t0\\\\\t5\\\\\\end{array} \\right] ,\\left[ \\begin{array}{c}\t4\\\\\t-5\\\\\t10\\\\\\end{array} \\right] \\,\\,belong\\,\\,to\\,\\,\\mathrm{Im}\\left( T \\right) . \\\\These\\,\\,vectors\\,\\,are\\,\\,linearly\\,\\,independent\\,\\,\\sin ce\\,\\,their\\,\\,coordinates\\,\\,are\\,\\,not\\,\\,proportional.\\\\Since\\,\\,dim\\left( \\mathrm{Im}\\left( T \\right) \\right) =2, \\left[ \\begin{array}{c}\t1\\\\\t0\\\\\t5\\\\\\end{array} \\right] ,\\left[ \\begin{array}{c}\t4\\\\\t-5\\\\\t10\\\\\\end{array} \\right] \\,\\,are\\,\\,the\\,\\,basis\\,\\,of\\,\\,\\mathrm{Im}\\left( T \\right)"